Pandora has hired 11 employees to make sales calls to prospective customers.

They each make 72 calls per day and have 61 people say that they will subscribe using the code you provide them. However, only a total of 60 actually subscribe using that code.

What is this group's process yield? (Keep three decimal places in your answer
Answer #1

Process yield is the ratio of successful calls to the calls attempted.

Successful calls = 60 since these are the only ones who actually subscribed.

Attempted calls = number of employees * number of calls made by each employee =11*72 = 792 calls are made

Of which ,only 60 were successful.

Group's process yield = 60 /792 = 0.015 or 1.150 %